# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ A.1 Notations
A.2 Lemmas for Proving Theorem 1
Lemma 1
@@ A.2 Lemmas for Proving Theorem 1
Proof 1
The length of epoch $m$ is given by $N_{m}=\lceil K\lambda_{m}2^{2(m-1)}\rceil\geq 2^{2(m-1)}$. From the lower bound of $N_{m}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{2^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=2^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq 2^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold.
Lemma 2
@@
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$.
Lemma 4
@@
Proof 5
The proof of Lemma 5 mainly follows the proof of Lemma 4 in Gupta et al. [2019]. We only consider the randomness in epoch $m$ and condition on all random variables before epoch $m$. In epoch $m$, the agent pulls arm $k$ with probability $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
@@
By applying the Chernoff-Hoeffding inequality (Theorem 1.1 in Dubhashi and Panconesi [2009], we can get
@@
By Freedman concentration inequality (Theorem 1 in Beygelzimer et al. [2011]), with probability at least $41-\beta_{m}/$ 4 we have
@@
Since $q_{k}^{m}=\widetilde{n}_{k}^{m}/N_{m}$, $\sum_{t\in E_{m}}|C_{k}^{t}|\leq C_{m}$ and $n_{k}^{m}\geq\lambda_{m}\geq 16\ln(4/\beta_{m})$, with probability at least $41-\beta_{m}/$ 4 the following inequality holds:
@@
Similarly, with probability at least $41-\beta_{m}/$ 4 the following inequality holds:
$-\frac{B_{k}^{m}}{\widetilde{n}_{k}^{m}}\leq\sqrt{\frac{\ln(4/\beta_{m})}{16\widetilde{n}_{k}^{m}}}+\frac{2C_{m}}{N_{m}}.$
@@
Combine Eq. (1) and Eq. (2), we complete the proof.
@@
Now we can we define the offset level $D_{m}$ for each epoch $m$:
@@
$-\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
@@
$-\frac{D_{m}}{N_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for the case of $m-1$, we have
@@
A.3 Proof for Theorem 1
@@
[…]
@@
$\begin{split}{\mathbb{E}}[R(T)]&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta)}{\Delta}+\log(8/\Delta)\\ &=O\left(C+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
Appendix B Proof of Theorem 2
B.1 Lemmas for Proving Theorem 2

### v2

@@ B.1.1 Notations
B.1.2 Lemmas for Proving Theorem 1
Lemma 1.
@@ B.1.2 Lemmas for Proving Theorem 1
Proof.
The length of epoch $m$ is given by $N_{m}=\lceil K\lambda_{m}2^{2(m-1)}\rceil\geq 2^{2(m-1)}$. From the lower bound of $N_{m}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{2^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=2^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq 2^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold. ∎
Lemma 2.
@@
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$. ∎
Lemma 4.
@@
Proof.
The proof of Lemma 5 mainly follows the proof of Lemma 4 in [gupta2019better]. We only consider the randomness in epoch $m$ and condition on all random variables before epoch $m$. In epoch $m$, the agent pulls arm $k$ with probability $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
@@
By applying the Chernoff-Hoeffding inequality (Theorem 1.1 in [dubhashi2009concentration]), we can get
@@
By Freedman concentration inequality (Theorem 1 in [beygelzimer2011contextual]), with probability at least $1-\beta_{m}/4$ we have
@@
Since $q_{k}^{m}=\widetilde{n}_{k}^{m}/N_{m}$, $\sum_{t\in E_{m}}|C_{k}^{t}|\leq C_{m}$ and $n_{k}^{m}\geq\lambda_{m}\geq 16\ln(4/\beta_{m})$, with probability at least $1-\beta_{m}/4$ the following inequality holds:
@@
Similarly, with probability at least $1-\beta_{m}/4$ the following inequality holds:
@@
Combine Eq. (2) and Eq. (3), we complete the proof. ∎
@@
Now we can define the offset level $D_{m}$ for each epoch $m$:
@@
$\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
@@
$\frac{D_{m}}{N_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for the case of m1m-1, we have
@@
where the third inequality comes from the upper bound of $\Delta_{k}^{m-1}$. ∎
B.1.3 Proof for Theorem 1
@@
We define ${\mathcal{A}}^{m}\triangleq\left\{k\in[K]\,\big|\,0<\Delta_{k}\leq 64\rho_{m-1}\right\}$ for epoch $m$. By combining all three cases, we can upper bound the regret as
@@
$\begin{split}R(T)&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta)}{\Delta}+\log(8/\Delta)\\ &=O\left(C+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
B.2 Proof of Theorem 2
B.2.1 Notations
B.2.2 Lemmas for Proving Theorem 2
Lemma 8.
For BARBAT with time horizon $T$, the number of epochs $M$ is at most $\log(VT)$. In the $m$-th epoch, the selected arm $k_{m}$ must satisfy $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ for all agents $v\in[V]$.
Proof.
The length of epoch $m$ is given by $N_{m}=\lceil K\lambda_{m}2^{2(m-1)}\rceil\geq 2^{2(m-1)}/V$. From the lower bound of $N_{m}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{2^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=2^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq 2^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold. ∎
Lemma 9.
For epoch $s$ and $m$ with $1\leq s\leq m$, the following inequality holds:
$\frac{\lambda_{m}}{\lambda_{s}}\leq\left(\frac{7}{6}\right)^{m-s}.$
Proof.
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$ and $V\geq 1$. ∎
Lemma 10.
For any epoch $m$, the following inequality holds:
$\sum_{s=1}^{m}\lambda_{s}\leq 2^{8}(m^{2}+m(10+\ln(VK))).$
Proof.
$=\sum_{s=1}^{m}2^{8}(2(s+4)\ln(2)+\ln(s+4)+\ln(4VK^{2}\ln(VK))$
$\leq\sum_{s=1}^{m}2^{8}(2s+7+\ln(4V^{2}K^{3}))$
$=2^{8}(m^{2}+8m+m\ln(4V^{2}K^{3}))$
$\leq 2^{8}(m^{2}+m(10+3\ln(VK))).$
Lemma 11.
Proof.
We can divide the sum to two components:
By applying the Chernoff-Hoeffding inequality (Theorem 1.1 in [dubhashi2009concentration]), we can get
By applying the concentration inequality for martingales (Theorem 1 in [beygelzimer2011contextual]), with probability at least $1-\frac{\beta_{m}}{4}$, we have
Since $q_{k}^{m}=\widetilde{n}_{k}^{m}/N_{m}$, $\sum_{t\in E_{m}}\sum_{v\in[V]}|C_{v,k}^{t}|\leq C_{m}$ and $n_{k}^{m}\geq\lambda_{m}\geq 16\ln(4/\beta_{m})/V$, with probability at least $1-\frac{\beta_{m}}{4}$, we can get the following inequality:
Similarly, $\frac{B_{k}^{m}}{V\widetilde{n}_{k}^{m}}$ also meets this bound with probability $1-\beta/4$. Therefore, we have
Combine Eq. 6 and Eq. 7, we complete the proof. ∎
We also define an event ${\mathcal{E}}_{m}$ for epoch $m$ as follows:
Lemma 12.
Proof.
By Lemma 11, we can get
A union bound over the $V$ agents and the $K$ arms conclude the proof.
Since $m\geq 1$, $V\geq 1$ and $K\geq 2$, then we can get
$VN_{m}$ $=KV2^{2(m-1)}\lambda_{m}$
$=K2^{2(m+3)}\ln((m+4)2^{2(m+5)}VK^{2}\ln(K))$
$\leq K2^{2(m+3)}(\ln(m+4)+2(m+5)\ln(2)+\ln(VK^{3}))$
$\leq K2^{2(m+4)}((m+4)\ln(VK))$
$\leq 1/\delta_{m}.$
$D_{m}=\begin{cases}2C_{m}&\textit{when ${\mathcal{E}}_{m}$ occurs}\\ VN_{m}&\textit{when ${\mathcal{E}}_{m}$ does not occur}\end{cases}.$
This way we can always guarantee the following inequality:
By the definition of $D_{m}$, we have
$\rho_{m}:=\sum_{s=1}^{m}\frac{D_{s}}{8^{m-s}VN_{s}}.$
Then we have the following lemma.
Lemma 13.
For all epochs $m$ and arms $k$, we can have
Proof.
$\frac{D_{m}}{VN_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{VN_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}.$
Additionally, since
we can get
$\frac{D_{m}}{VN_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{VN_{m}}.$
We now establish the upper bound for $\Delta_{k}^{m}$ by induction on $m$.
Assuming the statement is true for the case of m1m-1, we have
Where the second inequality follows from the induction hypothesis.
Next, we provide the lower bound of $\Delta_{k}^{m}$. We can get
where the third inequality comes from the upper bound of $\Delta_{k}^{m-1}$. ∎
B.2.3 Proof for Theorem 2
Then we analyze the regret in following three cases:
Case 1:
$0<\Delta_{k}\leq 64\rho_{m-1}$.
If $k\neq k_{m}$, then we have
$R_{v,k}^{m}=\Delta_{k}n_{k}^{m}\leq 64\rho_{m-1}n_{k}^{m}.$
If $k=k_{m}$, then we have
$R_{v,k}^{m}=\Delta_{k}\widetilde{n}_{k_{m}}^{m}\leq 64\rho_{m-1}N_{m}.$
Case 2:
$\Delta_{k}\leq 8\cdot 2^{-m}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
If $k\neq k_{m}$, then we have
If $k=k_{m}$, since $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$, we can get:
Case 3:
$\Delta_{k}>8\cdot 2^{-m}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
By Lemma 7 we have
We define ${\mathcal{A}}^{m}\triangleq\left\{k\in[K]\,\big|\,0<\Delta_{k}\leq 64\rho_{m-1}\right\}$ for epoch $m$. By combining all three cases, we can upper bound the regret for agent $v$ as
$\mathbb{E}[D_{m}]=2(1-\delta_{m})C_{m}+\delta_{m}N_{m}\leq 2C_{m}+1$
and we can bound $\sum_{m=1}^{M}\rho_{m-1}N_{m}$ as
Combining Eq. (8) and Eq. (9), by Lemma 10, we can get
$\begin{split}R_{v}(T)&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(VT)+3\log(VT)\log(30VK))}{V\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30VK)))}{V\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880\frac{C_{m}}{V}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(VT)\log(30VKT)}{V\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240VK/\Delta)}{V\Delta}+\log(8/\Delta)\\ &=O\left(\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log(T)\log(VKT)}{V\Delta_{k}}+\frac{K\log(1/\Delta)\log(VK/\Delta)}{V\Delta}\right).\end{split}$
So the cumulative regret for all agents as
$R(T)=O\left(C+\sum_{\Delta_{k}>0}\frac{\log(VT)\log(VKT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(VK/\Delta)}{\Delta}\right).$
$\textrm{Cost}(T)=\sum_{v\in[V]}M=O(V\log(VT)).$
B.3 Proof of Theorem 3
B.3.1 Notations
B.3.2 Lemmas for Proving Theorem 3
Lemma 14.
For the BB-BARBAT algorithm time horizon $T$, the number of epochs $M=L$. In the $m$-th epoch, the selected arm $k_{m}$ must satisfy $\Delta_{k_{m}}^{m-1}=a^{-(m-1)}$.
Proof.
The length of epoch $m$ is given by $N_{m}=K\lambda_{m}a^{2(m-1)}\geq a^{2(m-1)}$. From the lower bound of $N_{m}$ and $a=T^{\frac{1}{2(L+1)}}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{a^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=a^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq a^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold. ∎
Lemma 15.
For any epoch $m$, the length $N_{m}$ satisfies $N_{m}\geq\sum_{k\in[K]}n_{k}^{m}$.
Proof.
Lemma 16.
For epoch $s$ and $m$ with $1\leq s\leq m$, the following inequality holds:
$\frac{\lambda_{m}}{\lambda_{s}}\leq\left(\frac{7}{5}\right)^{m-s}.$
Proof.
We first show that the function $f(x)=\frac{x\ln(a^{2})+\ln(x)}{(7/5)^{x}}$ is strictly decreasing for $x\geq 5$, where $a=\max\{T^{\frac{1}{2(L+1)}},2\}\geq 2$. Notice that the derivative function
$<$ $\,\frac{\ln(m+4)+(m+4)\ln(a^{2})}{\ln(s+4)+(s+4)\ln(a^{2})}$
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$. ∎
Lemma 17.
For any epoch $m$, the following inequality holds:
$\sum_{s=1}^{m}\lambda_{s}\leq 2a^{8}\ln(a)(m^{2}+3m(2+\ln(K))).$
Proof.
Since $a\geq 2$ and $K\geq 2$, we can have
$=\sum_{s=1}^{m}a^{8}(2(s+4)\ln(a)+\ln(s+4)+\ln(4K^{2}\ln(aK))$
$\leq\sum_{s=1}^{m}a^{8}(2(s+4)\ln(a)+2(s+1)\ln(a)+6\ln(K)\ln(a))$
$\leq\sum_{s=1}^{m}a^{8}\ln(a)(4s+10+6\ln(K))$
$=2a^{8}\ln(a)(m^{2}+3m(2+\ln(K))).$
Lemma 18.
Proof.
By Lemma 5, we can get
A union bound over the $K$ arms conclude the proof.
Since $m\geq 1$, $K\geq 2$ and $a\geq 2$, then we can get
$N_{m}$ $=K\lambda_{m}a^{2(m-1)}$
$=Ka^{2(m+3)}\ln(4K^{2}(m+4)a^{2(m+4)}\ln(aK))$
$\leq Ka^{2(m+3)}(2(m+4)\ln(a)+\ln(m+4)+\ln(K^{2}\ln(aK))+\ln(4))$
$\leq Ka^{2(m+3)}(2(m+4)\ln(aK)+(m+4)\ln(aK)+3\ln(aK)+2\ln(aK))$
$\leq Ka^{2(m+3)}((4m+16)\ln(aK))$
$=Ka^{2(m+4)}((m+4)\ln(aK))$
$=1/\delta_{m}.$
This way we can always guarantee the following inequality:
$\rho_{m}:=\sum_{s=1}^{m}\frac{2^{m-s}D_{s}}{a^{4(m-s)}N_{s}}.$
Lemma 19.
For all epochs $m$ and arms $k$, we can have
Proof.
$\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
Additionally, since
we can get
$\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
We now establish the upper bound for $\Delta_{k}^{m}$ by induction on $m$.
Assuming the statement is true for the case of m1m-1, we have
Where the second inequality follows from the induction hypothesis.
Next, we provide the lower bound of $\Delta_{k}^{m}$. We can get
where the third inequality comes from the upper bound of $\Delta_{k}^{m-1}$. ∎
B.3.3 Proof for Theorem 3
We first define the regret $R_{k}^{m}$ generated by arm $k$ in epoch $m$ as
Then we analyze the regret in following three cases:
Case 1:
$0<\Delta_{k}\leq 64\rho_{m-1}$.
If $k\neq k_{m}$, then we have
$R_{k}^{m}=\Delta_{k}n_{k}^{m}\leq 64\rho_{m-1}n_{k}^{m}.$
If $k=k_{m}$, then we have
$R_{k}^{m}=\Delta_{k}\widetilde{n}_{k_{m}}^{m}\leq 64\rho_{m-1}N_{m}.$
Case 2:
$\Delta_{k}\leq 4\cdot a^{-(m-1)}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
If $k\neq k_{m}$, then we have
If $k=k_{m}$, since $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$, we can get:
Case 3:
$\Delta_{k}>4\cdot a^{-(m-1)}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
By Lemma 7 we have
We define ${\mathcal{A}}^{m}\triangleq\left\{k\in[K]\,\big|\,0<\Delta_{k}\leq 64\rho_{m-1}\right\}$ for epoch $m$. By combining all three cases, we can upper bound the regret as
$\mathbb{E}[D_{m}]=2(1-\delta_{m})C_{m}+\delta_{m}N_{m}\leq 2C_{m}+1$
and we can bound $\sum_{m=1}^{M}\rho_{m-1}N_{m}$ as
Combining Eq. (10) and Eq. (11), by Lemma 17, since $a=T^{\frac{1}{2(L+3)}}$ and $M=L$, we can get
$\begin{split}R(T)&\leq\sum_{m=1}^{M-1}570a^{2}{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &=\sum_{m=1}^{L-1}570a^{2}{\mathbb{E}}[D_{m}]+\sum_{m=1}^{L}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{L-1}1140a^{2}C_{m}+570a^{2}+2^{5}a^{8}\ln(a)\left(\sum_{\Delta_{k}>0}\frac{L^{2}+3L(2+\log(K))}{\Delta_{k}}\right.\\ &\quad\left.+\frac{K(\log^{2}(8/\Delta)+3\log(8/\Delta)(2+\log(K))}{\Delta}\right)+\log_{2}(8/\Delta)\\ &=O\left(CT^{\frac{1}{L+3}}+T^{\frac{4}{L+3}}\left(\sum_{\Delta_{k}>0}\frac{L\log(KT)}{\Delta_{k}}+\frac{K\log(T)\log(1/\Delta)\log(K/\Delta)}{L\Delta}\right)\right).\end{split}$
B.3.4 Proof for Theorem 4
Proof.
In the batched bandit problem, let the static grids be defined as $0=t_{0}<t_{1}<\cdots<t_{L}=T$. According to Lemma 2 in [gao2019batched], the lower bound for stochastic batched bandits with no corruption are given by:
$R(C=0,L)=O\left(K\cdot\max_{j\in[L]}\frac{t_{j}}{t_{j-1}+1}\right).$ (12)
For adversarial corruptions, we define an adversary who starts attacking at round 0 and continues until the end of a batch, denoted by $t_{s}$ for $0<s<L$. The attack strategy involves setting the rewards of all arms to zero, forcing any algorithm to pull all arms evenly in round $s+1$, which leads to a regret of approximately $t_{s}-t_{s-1}$. Therefore, under adversarial corruptions, the lower bound can be expressed as:
$\begin{split}R(C,L)&=O\left(K\cdot\max_{j\in[L]}\frac{t_{j}}{t_{j-1}+1}+C\cdot\max_{j\in[L-1]}\frac{t_{j+1}-t_{j}}{t_{j}}\right)\\ &=O\left(Kt_{1}+(K+C)\max_{j\in[L-1]}\frac{t_{j+1}-t_{j}}{t_{j}}\right)\\ &=O\left(Kt_{1}+(K+C)\left(\frac{T}{t_{j}}\right)^{\frac{1}{L-1}}\right).\end{split}$ (13)
According the formula, we can get the optimal value of $t_{1}=T^{\frac{1}{L}}\left(\frac{K(L-1)}{K+C}\right)^{\frac{L-1}{L}}$. Combine equation 13, we can get
$R(C,L)\geq\Omega\left(T^{\frac{1}{L}}\left(K+C^{1-\frac{1}{L}}\right)\right).$
B.4 Proof of Theorem 5
B.4.1 Lemmas for Proving Theorem 5